Computer science digital logics
Convert the following as specified DEC in hex to i
Answer #1

1(a)

DEC in hex to its decimal

(DEC)16 = ( )10

= D*162   +   E*161   + C*160

= 13*256 + 14*16 + 12*1

= 3328 + 224 +12

= (3564)10

b)

(DEC)16 to BCD

(DEC)16 = (3564)10

(3564)10 to BCD

3= 0011

5 = 0101

6 = 0110

4= 0100

BCD = 0011 0101 0110 0100

2(a)

1001 0111 0101 0010 BCD to octal

= 1001 =9, 0111=7, 0101=5, 0010 =2 from BCD to decimal =(9752)10

= (9752)10    to octal

                                             remainders

8|9752

8|1219                                     0

8|152                                       3

8|19                                         0

8|2                                           3

|0                                             2

taking remainders in reverse

=(23030)8

b)

(7643)8     to hex

7 = 111

6=110

4=100

3= 011

three bits for octal number

= 111 110 100 011

taking 4 bits to denote hexadecimal number

= 1111 1010 0011

= (FA3)     in hexadecimal

=(FA3)16